OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 16 and a15 = -40 ?

OpenStudy (cwrw238):

nth term of an AP = a1 + (n - 1)d

OpenStudy (anonymous):

Any general n-th term of an A.P. is , a_n = a_1 + (n-1)d Find 'd' and 'n' is known...

OpenStudy (cwrw238):

plug in a1 = 16, n = 15 15th term -41 to get value of common difference d

OpenStudy (cwrw238):

41st term = 16 + (41 - 1)d

OpenStudy (anonymous):

56(d)

OpenStudy (cwrw238):

no 16 + 40d

OpenStudy (cwrw238):

but you need to find d

OpenStudy (cwrw238):

sorry i had that wrong the 15th term is -40 so: -40 = 16 + (13 - 1) d now can you solve this for d?

OpenStudy (cwrw238):

oohh whats wrong with me n = 15 so its -40 = 16 + (15 - 1) d

OpenStudy (cwrw238):

can u do that?